A lake of area 200km^2 is covered by a slab of ice
Answer #1

Conductance through ice = kA/t

= 2.18*200*10^6/20 = 2.18*10^7 W/K

Current = 2.18e7*(20-4)

= 3.488*10^8 W

Thia current is too huge to drive a pick up